Problem 4-7 (add the formulas)

COMPOUND INTEREST AT NON-ANNUAL RATES

Given the annual interest rates below, and the compounding period indicated,

calculate the FUTURE VALUE for each item.

Annual

Compounding

Present Amount (SAR)

Interest

Years

Period

20,000

6%

5

Semi-annually

42,000

4%

12

Quarterly

8,000

12%

4

Monthly
Answer #1
Present Amount Number of times interest is compounded in a year Rate for a period Total Number of periods Future value = Present value(1+i)^n
20000 2 [2semiannual period in a year comprising of 6 months each] 6 /2 = 3% 5*2 =10

20000(1+.03)^10

20000(1.03)^10

20000*1.34392

26878.4   (rounded to 26878)

42000 4 [4quarters in a year comprising of 3 months each] 4/4 = 1% 12*4= 48

42000(1+.01)^48

42000* 1.61223

67713.66 (rounded to 67714)

8000 12 [12 months in a year] 12/12= 1% 4*12=48

8000(1+.01)^48

8000* 1.61223

12897.84 (rounded to 12898)